Which Graph Represents the Relationship x-2y=6


Answer by Theo(13342)

here's what the graph looks like.


y-intercept = -3
x-intercept = 6
slope = 1/2

all the different froms of the equation shown on the graph point to the same line on the graph.

that's because they're equivalent.

those equations and the common line are in red.

the lines in black intersect with the line in red. they are there to show the coordinate points.

you can see from the graph that the y-intercept is the point (0,-3) and the x-intercept is the point (6,0).

the y-intercept is the value of y when x = 0.
the x-intercept is the value of x when y = 0.

the slope is the change in the value of y divided by the corresponding change in the value of x.

you can choose any two points on the line to find the slope.
on this graph, the two points chosen are (0,-3) and (6,0).

you go from (0,-3) to (6,0).
slope is (0 - -3) / (6 - 0) = 3/6 = 1/2.

the x-axis is the vertical line where the value of x = 0.
the y-axis is the horizontal line where the value of y = 0.